. Define V =


 x


y





: x, y ≥ 0





. Decide if V is a vector space or not and prove


your claim. (Hint: V is the first quadrant in the xy-plane).

1
Expert's answer
2022-04-19T01:00:51-0400

"\left( 1,2 \right) ,\left( -2,-1 \right) \in V\\\left( 1,2 \right) +\left( -2,-1 \right) =\left( -1,1 \right) \notin V\\Not\,\,a\,\,vector\,\,space"